A Bibliography of Publications in the C++ Report

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

19 September 2010 Version 1.06

Title word cross-reference

-is [BN96c].

/Views [Hor91b].

7 [Hor92b].

Abstract [Mar92, Dew92, Jak97]. abstractions [Cog93]. access [Ped95]. action [Gla96a]. actively [Sch96]. ado [Cog91a]. Advantages [Whi96]. Algebra [BN95a]. algorithm [Tea93]. algorithms [Gla96a, Koe96b]. allocation [Han93]. also [Box96]. Alternative [Van96b, SV96c, SV96a, SV96b, TS95]. analysis [BN95b, BV91b, Mar94, Sch94a]. ANSI [BN96b, Sak91, Sak92]. ANSI/ISO [BN96b]. anyway [Dru91]. application [Hav94, KK95, Lan91]. applications [MNR97, Van96a, VS95]. approach [THN94]. Approaches [Mey92]. architecture [For96b, For96c, For96d, Kef93, KK95, MNR97]. Armor [Dav94]. array [Cog92a, Han93, Ped95]. arrays [Cog92a, SS93, SS94]. assignment [Koe96c]. association [Lin96b]. AT&T [Cop91]. AT&T’s [Mue96]. Automatic [BN96a]. automating [Mey96b]. Avoiding [Lak96b].

balance [Str93c]. base [Car92b]. based [Hor92a, KK95]. bases [Sak93]. be [Sak93]. begs [Shu97]. benefits [Ree93a]. Best [Cog91a, Laj96d]. better [LeJ91]. Beyond [Sou92a]. binary [Ham96, Lin96b]. bloat [Han97]. Booch [BV93]. Borland [Hor91b, Hor91a, Hor92a]. Bounding [Mey95]. Building [Box95a, KL97]. business [Fla95].

C [Ham96, Ham97, Ham92, Hor91b, Hor91a, Hor92b, Hor92a, KL96a, KL97, Lea93, Leg92, Ree96a, Sak91, Sak92, SLHS94, AOK94, ABB^{+}93, Bal92, Ban92, BR92, BN93, BN95a, BN96c, BN96d, BV91a, BV91c, Box95a, Bul96, BM92, Car96, Cog91b, Cog92c, Coh96a, Coh96b, Dav94, Dav97, Eck91, Fit96a, Fit96b, FR95, Han93, Hav94, Hic95, Hor91b, Hor95, Hor96a, Hun91, JS97, Kam96, KK95, Laj96b, Laj96a, Laj96c, Lak96a, Lak96c, Lak96b, Lan91, LeJ91, Lea96, Lin96b, Mar97a, Mar97b, Mey92, Mey93, Mue96, Mye93, Ree93a, Ree96c, Rei93, Sch92, Sch93a, Sch94b, Shu97, Sou92a, Sou92c, SD95, SS93, Ste91, Str93a, SL92, Str93b, Tav92, Tea93, TB97, Van96b, Vel95b, Vil95, VK93, Vin94, WD93, Wil92]. C/C [Hor92b]. Callbacks [Hic95, VS96]. can [LeJ91]. cards [Car91, Cop91]. case [Fit96a, Sch94b]. categories [BV91c]. Chain [OH96a]. cladding [Dav94]. class [Bec92, BV91c, Car92b, HD96, Hor92a, Hor96a, Kam96, Laj94a, Mus92, OH97, Ped95]. Classes [Shu97, CE95, Dav94, Mar92, Mot96, Sou96a]. client [Jak97]. closed [Mar96d]. Code [Cop96a, Box95b, CW97, Cog92c, Han97, Hor95]. Com [Box96]. Combining [LK97]. command [OH96a]. Committee [Laj96c]. communication [HD96, JS97, VS96]. comp.lang.C [Cog91a]. compare [Shu97]. Comparing [SV96c, SV96a, SV96b, TS95]. comparison [Mar97b, Shu97]. comparisons [Koe96a]. compile [Lak96b]. compile- [Lak96b]. compiler [Dru91, Hor92b]. component [BV92a, For96a]. Components [BV93, Leg92, Box95a, Car93]. computer [Rem95]. computing [Vin93]. Concurrency [AOK94]. Connector [Sch96]. consider [Str93c]. considered [Box96]. const [Mey96b]. constructors [Cog92b, Eck92]. Containers [Eck91, Koe96a, Han97]. controversy [Sou96a]. conventional [Cog92c]. conventions [Cog91b]. conversions [Laj95, Mey96b, Mey96d]. converting [BN96d, Fit96a, JS97]. cooperating [Sou96a]. Coping [Ree96b]. Copying [Koe96b]. CORBA [AG96, SV96a, SV96b, Vin93, Vin94, VS96]. CRC [Car91, Cop91]. critical [Mar97b]. Cross [Ban92, Jeg96]. Cross-over [Ban92]. Crossing [Tav92].

daemon [Sch94a]. data [OH96b, OH96c]. database [Str93a]. databases [BR92, Rob94, Sou92b, Str92]. dealing [Mar97a]. debugging [Ree96c]. Decentralized [SSH97]. decoupled [VS96]. deep [Str92]. delete [Laj96a]. delicate [Str93c]. Demeter [SLHS94]. dependencies [Lak96b]. Dependency [Mar96a]. depth [Hor91a]. Design [Car93, Mar97a, BV91a, BV91b, BV91d, BV92a, BV92b, BV92c, Cog93, Cop94, Hor95, Kef93, Lak96a, Lak96c, Lak96b, LK97, Mar94, Mar96b, Sch93a, Sch94b, Sch94a, SS95, Sch96, Str93b, Vli96a, BV91c]. design-C [BV91c]. Designing [Lin96a]. Developers [Ham92, Box95b]. Developing [CW97, For96a]. development [BV91e, Fla95, Hor96b, Jeg96, Mey93]. Dialog [OH96b]. differentiation [BN96a]. Dimensional [BN95b]. dimensions [Sch94a]. direction [Cop94]. DirectToSOM [Ham96, Ham97]. disadvantages [Whi96]. Distributed [Dew92, Vin93, VS96, TS95, VS95, Ham97]. do [LeJ91]. DOC [VS95]. document [For96b, For96c, For96d]. documents [BV91d]. does [Cog91c]. Doing [FR95, Dru91]. domain [Sch94a]. Don [Box95b]. done [Box95b]. DOS/C [Lan91]. driven [Sch93b]. dual [Mar97a]. Dynamic [Coh96a, OH96c, Car92a, Lin96a, SS93, SS94].

easily [Box95b]. Easing [Bec92]. Edition [Ham92]. editor [SS94]. Education [LeJ91]. Effective [Lak96c]. effectively [Mey92, Mey93]. effectiveness [Mey92]. efficient [Lin96a]. emerging [Cop94]. enable [Car92b]. Encapsulating [Sch93b, Sch92]. end [Leg92]. Engineering [Ree93a, BN93, BN96c, Mou94]. English [Box95b]. Enhancing [Han93]. environment [McC92]. equations [Van91]. Evaluating [HS96, HD96]. Event [Rob94, Sch93b, SSH97]. event-driven [Sch93b]. Events [SV97]. everywhere [WD93]. evolution [Sch94b]. evolve [SS95]. Examining [Mey93]. example [Mar94, VS95]. Exception [Chu94, Laj94b, Laj94a, Lea96, Mul96, Ree96d]. Exceptions [Ree93b, Ree96c, Ree96a, EC95, Ree96b, Ree97]. excessive [Lak96b]. exchange [OH96b]. execution [Cog92c]. Experience [Cop91]. Experiences [JS97, Lak96a]. explains [Box95b]. Exploring [Shu97]. Expression [Vel95a]. Extending [For96b, For96c, Hor94, Mus92]. extensions [Str93c].

factories [OH97]. failed [Cog92b]. feature [Ree93b]. features [Van96a]. finalized [Sak93]. Fine [Mye93]. Fine-tuning [Mye93]. first [Hor91b, Hor92b, Laj94a, Rei93]. first-class [Laj94a]. fit [Str94]. formatting [Bec92]. four [Vli95]. Framework [OH97, BR95, JS97]. frameworks [Lan91]. friends [Van91]. Function [Laj96d, Dru91, Han93]. functions [Mar92]. functors [Hic95].

gamble [Mue96]. gang [Vli95]. Gap [Vli96a]. generated [Rem95]. Generating [BN97]. generation [Hor95, Vli96a]. Generative [Cop94, THN94]. Generic [Gla97, SS93, LK97, THN94]. Getting [Ske92a]. GISMO [ABB^{+}93]. GNU [Lea93]. gracefully [Cog92b]. grafting [BN93]. Granularity [Mar96b]. graphics [Fit96b, Lak96a]. groups [Sou96a].

Guarded [Sim97]. GUI [Jeg96, MNR97]. guidelines [Ree96d].

handle [Ske92a]. handler [Ske92a].

Handling [Cog92b, Laj94b, Laj94a, Lea96, Mul96, Rob94, SSH97]. harder [Car92a]. harmful [Box96]. hashing [Lin96a]. hatching [Vli95]. hatching-perspectives [Vli95]. Hello [BN96b]. Helper [Gla96a]. hierarchies [Mar97a]. hierarchy [Lak96c]. Hollywood [Vli96b]. human [Cop96b].

I/O [BR95, Sch93a, Sch93b]. identification [SL92]. Identifying [Jor95]. IDL [Vin94]. imagery [Rem95]. implementation [Sch93a]. Implementing [Lin96a, Lin96b, Sou97, SS94, Bal92]. important [Box95b]. impression [Hor92b]. improved [Lak96c]. in-depth [Hor91a]. independent [Lak96c]. Inheritance [Koe96c, Car92b, Coh96a, Mar97a, Mey96b]. initialized [Sak93]. initializing [Sch96]. Inner [Shu97]. input [BN96c]. insertion [Koe96b]. instantiation [McC92]. instantiations [Mey95]. insulation [Lak96b]. integrating [SS93]. Interface [BL91, CE95, Sch93b]. interfaces [Dew92, Sch92]. interprocess [HD96]. Introducing [van94]. introduction [Vil94]. Invasive [Car92b]. Inversion [Mar96a]. iostream [Hor94]. IPC [Sch92]. IPCs [Sch93b]. ISO [BN96b]. issues [Ree93b]. iterator [KL97]. Iterators [Bec93, KL96a, KL96b, BN96d, Eck91].

Java [Dav97, Mar97b, Car96, Gla97, Hor96b, JS97, Shu97]. JDK [Shu97]. job [Bec92]. journey [Tav92].

key [Cog93]. kill [Vli96d]. kit [Hor96b].

Klingon [Box95b].

land [Ske92b]. landscape [Rem95].

language [Box95b, Mue96, Ree93b, Rei93, Str93c, Jor96]. languages [Box95b, Cop94, Cop97]. Large [Lak96a, Lak96c, Lak96b, Mar97c, Mue96]. Large-scale [Lak96a, Lak96c, Lak96b, Mar97c]. laundering [Vli97]. lessons [Jam96]. levelization [Lak96b]. libraries [BV92a, BV92c, Cog91b, HD96, Hor96a]. Library [Str93b, BV91a, Cog93, Hor92a, Hor94, Kam96, Lea93, Sou92c, TB97, Vil95, BL91, Gla97, KL96a, KL96b, KL97, Vil94]. Lightweight [Coh96b, Mot96]. link [Lak96b]. link-time [Lak96b]. Linking [Hun91]. Liskov [Mar96c]. LISP [Tea93]. Logical [BV92a]. logstrbuf [Mus92]. long [Cog91c]. look [Hor91b, Hor91a, Str92]. looks [Car92a].

maintainability [Lak96c]. Maintaining [Str93c]. maintenance [Mar96d]. Making [Str94, Ban92]. management [BV92b, BV92c, Fla95, Mye93]. manager [Ree93a]. managing [Kam96, Sou96a]. manipulations [Laj96b]. manipulators [Mou94]. mannered [Dav97]. Mapping [Vin94]. matures [Shu97]. mechanism [Laj94b]. members [Dew93]. membership [Vli96c]. Memory [Mye93, Sou92b, Han93, Kam96, Laj96b, Laj96a, Laj96e]. Memory-resident [Sou92b]. Mentor [Lak96a]. metaprograms [Vel95b]. methods [Van96b]. MFC [For96a, For96b, For96c, For96d, OH96b, OH96c, Oni96, Oni97]. Microsoft [Hor92b]. model [Laj96b, Laj96a, Laj96e]. model-type [Laj96e]. modelling [VS95]. modifying [Car92b]. moving [Sim97]. msgstream [Bec92]. much [Cog91a]. multimedia [Van96a]. Multiplatform [Jam96]. multiplexing [Sch93a, Sch93b]. multithreaded [Lea96, SV96c, SV96a, SV96b, Sim97, TB97]. Multithreading [Oni97]. myth [Bul96].

Naming [Cog91b]. nested [Shu97]. network [HS96, Sch94a, Sch96]. new-handler [Ske92a]. NT [SS95]. null [Cog91a]. nullness [Mey96d]. numerical [Hor96a].

O [BR95, Sch93a, Sch93b]. Object [BV91b, BV91c, BV91d, BV92b, Oni96, Str92, BR92, BR95, BV91e, Hor92a, LK97, Lip96, Mar94, Mar96b, Mey95, Rob94, Sch92, Sch93a, Sch93b, Str93a, Vin93, VS95, BL91, Jor96]. object-based [Hor92a]. Object-oriented [BV91b, BV91c, BV91d, BV92b, BR92, BR95, BV91e, LK97, Mar94, Mar96b, Rob94, Sch92, Sch93a, Sch93b]. objects [BR92, Dav97, Ham96, Jor95, Kam96, Laj94a]. Observer [Vis96]. ODMG [Jor95, Str93a]. ODMG-93 [Str93a]. old [Van96b]. OLE2 [Box95a]. OMG [SV97]. onto [BN93]. OO [Car91, HS96, Hor95]. OOP [BV92c, Mus92, Ree93a, Ske92b]. open [Mar96d]. open-closed [Mar96d]. operating [Sch93b]. Operator [Ped95, Wil92]. operators [BN95a, Laj96a]. OQL [Jor96]. Oracle [BM92]. organization [Laj96c]. oriented [BR92, BR95, BV91b, BV91c, BV91d, BV91e, BV92b, LK97, Mar94, Mar96b, Rob94, Sch92, Sch93a, Sch93b]. other [Box95b]. our [BN93]. Overcoming [Mue96]. overhead [Dru91]. overload [Laj96d]. overloading [Wil92]. overview [AOK94, Car96].

package [Leg92]. packages [Lak96b]. paradigms [Tav92]. parameterization [Cra94]. Pattern [Cop97, Vli95, Cop94, MNR97, Sch96, Vli96a]. Patterns [Jak97, Cop96a, Cop96b, Mar97a, SS95, Sou96b, Sou97, Whi96]. Performance [Bul96, HS96]. persistence [Coh96b, Oni96, THN94]. Persistent [BR92]. perspective [Ree93a]. perspectives [Vli95]. Physical [BV92c, BV92b, Lak96c]. pilgrim [Tav92]. platform [Jeg96]. Pointer [Laj96b, Koe96a]. pointers [Box96, Chu94, Mey96a, Mey96c, Mey96b, Mey96d, Sim97]. portability [Bec93]. portable [TB97]. Porting [Box95b, Dav97]. power [Van96a]. Practical [Cog93, Mar94]. practice [Jak97]. pragmatic [Ree93b]. primary [Mue96]. principle [Mar96c, Mar96d, Vli96b, Mar96a]. principles [Cog93]. privileges [Vli96c]. problems [Mue96, Van96b]. process [BV91e, Box95b, CW97]. productivity [Jam96]. program [Cog91c, Fit96b]. programmers [Car96]. Programming [For96d, BN93, Car91, HS96, Hor91b, Hor96a, Hor96b, LK97, Mue96, Rei93, Sch92, SV96c, SV96a, SV96b, TB97, TS95]. programs [Fit96a, Lea96, Mye93]. progress [Sak91, Sak92]. project [ABB^{+}93, FR95, Mue96]. Protection [Vli96b, Vli96c]. pure [Mar92].

Quality [Sou96b]. Query [Jor96]. Questioning [Cog92c]. questions [Cog92a].

Reactor [Sch93a]. real [CW97]. reality [Bul96]. realizing [Ree93a]. reallocation [Han93]. reallocation/renew [Han93]. recovery [Chu94]. recursive [Tea93]. Redesigning [Fit96b]. reduce [Han97]. Refinements [Mey96a]. Reflections [Ree97]. release [Leg92]. renew [Han93]. report [SLHS94]. representation [Whi96]. representations [Laj96e]. resident [Sou92b]. resolution [Laj96d]. Responsibility [OH96a]. reusable [CW97, Lin96b]. reuse [Bec93, FR95, Lak96c]. Reusing [Ham96]. Reverse [Mou94]. Reverse-engineering [Mou94]. revisited [Gau96]. right [Cog93]. Risky [Fla95]. robustness [Shu97]. roots [BN93]. rules [Mul96]. run [Cog91c]. Runtime [Cra94, SL92, Laj94b, VK93].

Saber [Ste91]. Saber-C [Ste91]. safe [Lin96a]. scale [Lak96a, Lak96c, Lak96b, Mar97c, Mue96]. schedule [Laj96c]. Scientific [BN93, BN96c]. search [Lin96a]. See [Shu97]. Selecting [Cog93, Sou92c]. selection [Laj96d]. separating [Bul96]. serialization [OH96c]. servers [SV96c, SV96a, SV96b]. service [AG96, Gla96b, SV97]. services [Sch92, Sch96]. shared [Kam96]. ShMem [Kam96]. should [Sak93]. side [Cop96b]. Simplifying [BV93]. Simultaneous [Van91]. singleton [Vli96d]. sixth [Laj96c]. size [Cog92c]. slicing [Koe96c]. Smalltalk [VK93]. Smart [Mey96c, Mey96b, Mey96d, Bec92, Box96, Chu94, Mey96a, Sim97]. software [BV92a, Cop94, Fla95, FR95, JS97, Lak96a, Lak96c, Lak96b, Mar96d, SS95, Vli96a]. solving [Van96b]. SOM [Ham97]. source [Box95b]. specification [Ree96d]. speed [Cog92c]. stability [Mar97c]. Standard [Car93, Leg92, BN96c, Str93a, Str94, Vil95, KL96a, KL96b, KL97, Ree96a, Vil94]. Standards [Laj95, Laj96c]. STL [Gla96a, Han97, Koe96b, KL97]. storage [BV92b, BV92c]. story [Hav94]. strategies [Bal92]. streambufs [Mus92]. streaming [AG96, Gla96b]. Strengthen [Van96a]. structure [OH96c]. structures [Lin96a]. study [Fit96a, Sch94b]. Subobject [Dew93]. substitution [Mar96c]. Subtleties [Wil92]. successfully [Mul96]. Sun [Hor96b]. support [AOK94, VK93]. supporting [Laj94b, Laj94a]. survey [Rem95]. system [Sch93b, SS95, van94]. Systems [Sch92, Lak96a, Sim97].

targets [OH96a]. Taskmaster [MNR97]. technique [Mye95]. techniques [CW97, SV96c, SV96a, SV96b, TS95]. Template [KL96b, Vil94, For96b, For96c, For96d, Hic95, KK95, McC92, Vel95b]. template-based [KK95]. templates [Bal92, Sou92a, SS93, Vel95a]. Ten [Mul96, Ree96d]. test [Fit96b]. testing [Lak96c]. text [Box95b, SS94]. Threads.h [TB97]. time [Lak96b]. tips [Dav97]. toolkits [HS96, Jeg96]. tools [Hor95, Mey93]. Tools.h [Kef93]. Tradeoffs [CE95, EC95]. training [Ban92]. traits [Mye95]. transforming [BN97]. transition [Ban92, Ree93a]. translator [Bec92]. transparency [Str92]. Transplanting [Tea93]. tree [Tea93]. tree-recursive [Tea93]. trouble [Vis96]. tuning [Mye93]. Type [Vli97, Laj95, Laj96e, Mey96d, SL92, VK93].

Unicode [Box95b]. unique [Whi96]. universal [AG96, Gla96b]. UNIX [Sch93a, Sch93b, SS95]. update [Laj95, Laj96c, Sak91, Sak92]. use [BN96d]. useful [Mye95]. user [Leg92]. Using [AG96, Ham97, Mey92, Mey93, SS95, Vel95b, Box95a, Fit96b, Hic95, Lak96a, SS94, Str93b]. USL [Car93, Leg92].

varieties [Lip96]. vector [Car92a, Str94]. Version [Hor92b, Ste91]. viable [Laj96d]. view [For96a]. Views [Hor91b]. Virtual [Eck92, Dru91, Mar92, Sak93]. Visitors [Gau96].

Web [SD95]. Well [Dav97]. Well-mannered [Dav97]. While [BN96d]. Wide [SD95]. Window [van94]. Windows [Box95b, Hor91b, SS95]. wisdom [Cog92c]. works [Sak91, Sak92]. world [CW97, SD95]. wrapper [Sch93a]. wrappers [Sch92]. written [Box95b].

X [van94].

year [Laj96c]. year-organization [Laj96c].

Zortech [Ham92].

References

Atwood:1993:GPC

[ABB^{+}93] W. Atwood, A. Breakstone, D. Britton, T. Burnett, D. Myers, and G. Word. The GISMO project (C++). C++ Report, 5(3): 38–43, March–April 1993. CODEN CRPTE7. ISSN 1040-6042.

Alicea:1996:UUS

[AG96] H. Alicea and G. Glass. Using the universal streaming service with CORBA. C++ Report, 8(8):12–19, September 1996. CODEN CRPTE7. ISSN 1040-6042.

Arjomandi:1994:CSC

[AOK94] E. Arjomandi, W. O’Farrell, and I. Kalas. Concurrency support for C++: an overview. C++ Report, 6 (1):44–50, January 1994. CODEN CRPTE7. ISSN 1040-6042.

Ball:1992:ITI

[Bal92] M. Ball. Inside templates: implementing C++ strategies. C++ Report, 4(7):36–40, September 1992. CODEN CRPTE7. ISSN 1040-6042.

Banahan:1992:CTM

[Ban92] M. Banahan. Cross-over training: making the transition from C to C++. C++ Report, 4(8):44–48, October 1992. CODEN CRPTE7. ISSN 1040-6042.

Becker:1992:MCS

[Bec92] P. Becker. A msgstream class for smart formatting: Easing the translator’s job. C++ Report, 4(7):42–48, September 1992. CODEN CRPTE7. ISSN 1040-6042.

Becker:1993:IPR

[Bec93] P. Becker. Iterators, portability, and reuse. C++ Report, 5(2):51–54, February 1993. CODEN CRPTE7. ISSN 1040-6042.

Brazille:1991:OIL

[BL91] R. Brazille and B. Leggett. The Object Interface Library. C++ Report, 3(7):18–20, 22, 24, July– August 1991. CODEN CRPTE7. ISSN 1040-6042.

Buroff:1992:CO

[BM92] S. Buroff and R. Murray. The C++ oracle. C++ Report, 4(4):9–10, May 1992. CODEN CRPTE7. ISSN 1040-6042.

Barton:1993:SEP

[BN93] J. J. Barton and L. R. Nackman.

Scientific and engineering programming in C++: grafting onto our roots. C++ Report, 5(7):26–31, September 1993. CODEN CRPTE7. ISSN 1040-6042.

Barton:1995:ACO

[BN95a] J. J. Barton and L. R. Nackman. Algebra for C++ operators. C++ Report, 7(3):70–74, March–April 1995. CODEN CRPTE7. ISSN 1040-6042.

Barton:1995:DA

[BN95b] J. J. Barton and L. R. Nackman.

Dimensional analysis. C++ Report, 7(1):39–40, 42–43, January 1995. CODEN CRPTE7. ISSN 1040-6042.

Barton:1996:AD

[BN96a] J. J. Barton and L. R. Nackman.

Automatic differentiation. C++ Report, 8(2):61–63, February 1996. CODEN CRPTE7. ISSN 1040-6042.

Barton:1996:HAI

[BN96b] J. J. Barton and L. R. Nackman. Hello ANSI/ISO. C++ Report, 8(4): 68–70, 72–73, April 1996. CODEN CRPTE7. ISSN 1040-6042.

Barton:1996:SEC

[BN96c] J. J. Barton and L. R. Nackman. Scientific and engineering C++-is this standard input? C++ Report, 8(9):66, 68–69, October 1996. CO-DEN CRPTE7. ISSN 1040-6042.

Barton:1996:WCC

[BN96d] J. J. Barton and L. R. Nackman. While (converting to C++) use iterators. C++ Report, 8(6):77–78, 83, 87, June 1996. CODEN CRPTE7. ISSN 1040-6042.

Barton:1997:GT

[BN97] J. J. Barton and L. R. Nackman. Generating and transforming. C++ Report, 9(2):62–64, February 1997. CODEN CRPTE7. ISSN 1040-6042.

Box:1995:BCC

[Box95a] D. Box. Building C++ components using OLE2. C++ Report, 7(3): 28–34, March–April 1995. CODEN CRPTE7. ISSN 1040-6042.

Box:1995:PKP

[Box95b] Don Box. Porting to Klingon — porting source code written to process English language text to other languages is not easily done. Don explains why Unicode is important for Windows developers. C++ Report, 7(4):15–19, May 1995. ISSN 1040-6042.

Box:1996:CSP

[Box96] D. Box. Com smart pointers also considered harmful. C++ Report, 8(3):54–57, March 1996. CODEN CRPTE7. ISSN 1040-6042.

Bartels:1992:POO

[BR92] D. Bartels and J. Robie. Persistent objects and object-oriented databases for C++. C++ Report, 4(7):49–50, 52–56, September 1992. CODEN CRPTE7. ISSN 1040-6042.

Berg:1995:OFI

[BR95] B. Berg and E. Rowlance. An object-oriented framework for I/O. C++ Report, 7(4):28–35, May 1995. CODEN CRPTE7. ISSN 1040-6042.

Bulka:1996:PCS

[Bul96] D. Bulka. Performance and C++:

separating myth from reality. C++ Report, 8(8):46–54, September 1996. CODEN CRPTE7. ISSN 1040-6042.

Booch:1991:CLD

[BV91a] G. Booch and M. Vilot. C++ library design. C++ Report, 3(6):10– 14, June 1991. CODEN CRPTE7. ISSN 1040-6042.

Booch:1991:OAD

[BV91b] G. Booch and M. Vilot. Objectoriented analysis and design. C++ Report, 3(8):7–10, September 1991. CODEN CRPTE7. ISSN 1040-6042.

Booch:1991:ODC

[BV91c] G. Booch and M. Vilot. Objectoriented design-C++ class categories. C++ Report, 3(7):6– 10, July–August 1991. CODEN CRPTE7. ISSN 1040-6042.

Booch:1991:ODD

[BV91d] G. Booch and M. Vilot. Objectoriented design documents. C++ Report, 3(10):8, 10–12, November– December 1991. CODEN CRPTE7. ISSN 1040-6042.

Booch:1991:ODP

[BV91e] G. Booch and M. Vilot. The objectoriented development process. C++ Report, 3(9):8–11, October 1991. CODEN CRPTE7. ISSN 1040-6042.

Booch:1992:LDS

[BV92a] G. Booch and M. Vilot. Logical design: software component libraries. C++ Report, 4(4):18, 20–22, May 1992. CODEN CRPTE7. ISSN 1040-6042.

Booch:1992:ODP

[BV92b] G. Booch and M. Vilot. Objectoriented design: physical design: storage management. C++ Report, 4(1):7–10, January 1992. CODEN CRPTE7. ISSN 1040-6042.

Booch:1992:PDS

[BV92c] G. Booch and M. Vilot. Physical design: storage and management libraries (OOP). C++ Report, 4 (2):8–10, February 1992. CODEN CRPTE7. ISSN 1040-6042.

Booch:1993:SBC

[BV93] G. Booch and M. Vilot. Simplifying the Booch components. C++ Report, 5(5):41–52, June 1993. CO-DEN CRPTE7. ISSN 1040-6042.

Cargill:1991:CCO

[Car91] T. Cargill. CRC cards (OO programming). C++ Report, 3(8):

11–13, September 1991. CODEN CRPTE7. ISSN 1040-6042.

Cargill:1992:DVH

[Car92a] T. Cargill. A dynamic vector is harder than it looks. C++ Report, 4(5):47–50, June 1992. CO-DEN CRPTE7. ISSN 1040-6042.

Carroll:1992:IIM

[Car92b] M. Carroll. Invasive inheritance:

modifying a base class to enable inheritance. C++ Report, 4(8):34–42, October 1992. CODEN CRPTE7. ISSN 1040-6042.

Carroll:1993:DUS

[Car93] M. Carroll. Design of the USL Standard Components. C++ Report, 5 (5):34–39, 53, June 1993. CODEN CRPTE7. ISSN 1040-6042.

Cargill:1996:OJC

[Car96] T. Cargill. An overview of Java for C++ programmers. C++ Report, 8 (2):46–49, February 1996. CODEN CRPTE7. ISSN 1040-6042.

Carroll:1995:TIC

[CE95] M. Carroll and M. Ellis. Tradeoffs of interface classes. C++ Report, 7 (1):29–32, January 1995. CODEN CRPTE7. ISSN 1040-6042.

Churchill:1994:ERS

[Chu94] S. Churchill. Exception recovery with smart pointers. C++ Report, 6(1):34–35, 38–42, January 1994. CODEN CRPTE7. ISSN 1040-6042.

Coggins:1991:BCA

[Cog91a] J. M. Coggins. Best of comp.lang.

c++ — much ado about null. C++ Report, 3(10):21–24, November– December 1991. CODEN CRPTE7. ISSN 1040-6042.

Coggins:1991:NCC

[Cog91b] J. M. Coggins. Naming conventions in C++ libraries. C++ Report, 3 (9):17–21, October 1991. CODEN CRPTE7. ISSN 1040-6042.

Coggins:1991:WDP

[Cog91c] J. M. Coggins. Why does this program run so long? C++ Report, 3(6):21–24, June 1991. CODEN CRPTE7. ISSN 1040-6042.

Coggins:1992:AQA

[Cog92a] J. M. Coggins. An array of questions about arrays. C++ Report, 4 (2):19–22, February 1992. CODEN CRPTE7. ISSN 1040-6042.

Coggins:1992:HFC

[Cog92b] J. M. Coggins. Handling failed constructors gracefully. C++ Report, 4 (1):20–22, January 1992. CODEN CRPTE7. ISSN 1040-6042.

Coggins:1992:QCW

[Cog92c] J. M. Coggins. Questioning conventional ‘wisdom’ about C++ code size and execution speed. C++ Report, 4(4):58–61, May 1992. CO-DEN CRPTE7. ISSN 1040-6042.

Coggins:1993:PPL

[Cog93] J. M. Coggins. Practical principles for library design. I. selecting the right abstractions is the key. C++ Report, 5(2):38–42, February 1993. CODEN CRPTE7. ISSN 1040-6042.

Cohen:1996:DIC

[Coh96a] S. Cohen. Dynamic inheritance in C++. C++ Report, 8(4):30– 37, April 1996. CODEN CRPTE7. ISSN 1040-6042.

Cohen:1996:LPC

[Coh96b] S. Cohen. Lightweight persistence in C++. C++ Report, 8(5):22–32, May 1996. CODEN CRPTE7. ISSN 1040-6042.

Coplien:1991:ECC

[Cop91] J. Coplien. Experience with CRC cards in AT&T. C++ Report, 3(8): 1, 4–6, September 1991. CODEN CRPTE7. ISSN 1040-6042.

Coplien:1994:GPL

[Cop94] J. O. Coplien. Generative pattern languages: an emerging direction of software design. C++ Report, 6 (6):18–22, 66–67, July–August 1994. CODEN CRPTE7. ISSN 1040-6042.

Coplien:1996:CP

[Cop96a] J. O. Coplien. Code patterns. C++ Report, 8(9):18–25, October 1996. CODEN CRPTE7. ISSN 1040-6042.

Coplien:1996:HSP

[Cop96b] J. O. Coplien. The human side of patterns. C++ Report, 8(1):73–74, 76–78, 80, January 1996. CODEN CRPTE7. ISSN 1040-6042.

Coplien:1997:PL

[Cop97] J. O. Coplien. Pattern languages.

C++ Report, 9(1):15–21, January 1997. CODEN CRPTE7. ISSN 1040-6042.

Crawford:1994:RP

[Cra94] J. Crawford. Runtime parameterization. 1. C++ Report, 6(7):24–29, September 1994. CODEN CRPTE7. ISSN 1040-6042.

Cochran:1997:DRC

[CW97] G. Cochran and R. Weisz. Developing reusable code in the real world: process and techniques. C++ Report, 9(4):22–31, April 1997. CO-DEN CRPTE7. ISSN 1040-6042.

Davis:1994:ACC

[Dav94] S. R. Davis. Armor cladding C++ classes. C++ Report, 6(8):36–39, 41, October 1994. CODEN CRPTE7. ISSN 1040-6042.

Davis:1997:WMO

[Dav97] M. Davis. Well-mannered objects: tips on porting C++ to java. C++ Report, 9(4):14–21, April 1997. CO-DEN CRPTE7. ISSN 1040-6042.

Dewhurst:1992:DAI

[Dew92] S. C. Dewhurst. Distributed abstract interfaces. C++ Report, 4(3): 44–50, March–April 1992. CODEN CRPTE7. ISSN 1040-6042.

Dewhurst:1993:SM

[Dew93] S. C. Dewhurst. Subobject members. C++ Report, 5(3):51– 53, March–April 1993. CODEN CRPTE7. ISSN 1040-6042.

Druker:1991:WCD

[Dru91] S. Druker. ‘what’s that compiler doing, anyway?’ — virtual function overhead. C++ Report, 3(6):19–20, June 1991. CODEN CRPTE7. ISSN 1040-6042.

Ellis:1995:TE

[EC95] M. Ellis and M. Carroll. Tradeoffs of exceptions. C++ Report, 7(3):12, 14, 16, March–April 1995. CODEN CRPTE7. ISSN 1040-6042.

Eckel:1991:CIC

[Eck91] B. Eckel. Containers and iterators in C++. C++ Report, 3(7): 12–13, July–August 1991. CODEN CRPTE7. ISSN 1040-6042.

Eckel:1992:VC

[Eck92] B. Eckel. Virtual constructors. 2.

C++ Report, 4(4):13–16, May 1992. CODEN CRPTE7. ISSN 1040-6042.

Fitzpatrick:1996:CSC

[Fit96a] J. Fitzpatrick. A case study: converting C programs to C++. C++ Report, 8(2):40–45, 63, February 1996. CODEN CRPTE7. ISSN 1040-6042.

Fitzpatrick:1996:RGT

[Fit96b] J. Fitzpatrick. Redesigning a graphics test program using C++. C++ Report, 8(6):54–60, June 1996. CO-DEN CRPTE7. ISSN 1040-6042.

Flanagan:1995:RBS

[Fla95] E. B. Flanagan. Risky business (software development management). C++ Report, 7(3):75–76, 78, March–April 1995. CODEN CRPTE7. ISSN 1040-6042.

Forguson:1996:DVC

[For96a] D. Forguson. Developing the view component in MFC. C++ Report, 8 (6):67–74, 76, June 1996. CODEN CRPTE7. ISSN 1040-6042.

Forguson:1996:EMDa

[For96b] D. Forguson. Extending the MFC document template architecture. C++ Report, 8(4):43–53, April 1996. CODEN CRPTE7. ISSN 1040-6042.

Forguson:1996:EMDb

[For96c] D. Forguson. Extending the MFC document template architecture. C++ Report, 8(9):57–65, October 1996. CODEN CRPTE7. ISSN 1040-6042.

Forguson:1996:PMD

[For96d] D. Forguson. Programming with the MFC document template architecture. C++ Report, 8(3):36–41, March 1996. CODEN CRPTE7. ISSN 1040-6042.

Flanagin:1995:DIC

[FR95] G. Flanagin and N. Rimmer. Doing it (C++ software reuse project). C++ Report, 7(1):24–28, January 1995. CODEN CRPTE7. ISSN 1040-6042.

Gautier:1996:VR

[Gau96] P. Gautier. Visitors revisited. C++ Report, 8(8):36–39, 42–45, September 1996. CODEN CRPTE7. ISSN 1040-6042.

Glass:1996:SAH

[Gla96a] G. Glass. STL in action. helper algorithms. C++ Report, 8(1):18–21, January 1996. CODEN CRPTE7. ISSN 1040-6042.

Glass:1996:USS

[Gla96b] G. Glass. A universal streaming service. C++ Report, 8(4):74–76, 81– 83, April 1996. CODEN CRPTE7. ISSN 1040-6042.

Glass:1997:JGL

[Gla97] G. Glass. The Java Generic Library. C++ Report, 9(1):70–74, January 1997. CODEN CRPTE7. ISSN 1040-6042.

Hamsath:1992:ZCD

[Ham92] N. Hamsath. Zortech C++ Developers Edition. C++ Report, 4 (2):23–25, February 1992. CODEN CRPTE7. ISSN 1040-6042.

Hamilton:1996:RBO

[Ham96] J. Hamilton. Reusing binary objects with DirectToSOM C++. C++ Report, 8(3):42–51, March 1996. CO-DEN CRPTE7. ISSN 1040-6042.

Hamilton:1997:UDS

[Ham97] J. Hamilton. Using Distributed SOM with DirectToSOM C++. C++ Report, 9(2):28–36, February 1997. CODEN CRPTE7. ISSN 1040-6042.

Hansen:1993:EMA

[Han93] T. L. Hansen. Enhancing memory allocation. an array reallocation/renew function for C++. C++ Report, 5(6):20–22, 24, 26–30, July– August 1993. CODEN CRPTE7. ISSN 1040-6042.

Hansen:1997:HRC

[Han97] M. Hansen. How to reduce code bloat from STL containers. C++ Report, 9(1):34–41, January 1997. CODEN CRPTE7. ISSN 1040-6042.

Havener:1994:CAS

[Hav94] C. L. Havener. A C++ application story. C++ Report, 6(2):31–33, February 1994. CODEN CRPTE7. ISSN 1040-6042.

Henkel:1996:ECL

[HD96] S. Henkel and W. Dangelmaier.

Evaluating class libraries for interprocess communication. C++ Report, 8(7):21–29, July 1996. CODEN CRPTE7. ISSN 1040-6042.

Hickey:1995:CCU

[Hic95] R. Hickey. Callbacks in C++ using template functors. C++ Report, 7 (2):42–50, February 1995. CODEN CRPTE7. ISSN 1040-6042.

Horstmann:1991:ILB

[Hor91a] C. Horstmann. An in-depth look as Borland C++. C++ Report, 3(10): 17–20, November–December 1991. CODEN CRPTE7. ISSN 1040-6042.

Horstmann:1991:FLA

[Hor91b] C. S. Horstmann. A first look at programming Windows with C++/ Views and Borland C++. C++ Report, 3(6):15–18, 24, June 1991. CO-DEN CRPTE7. ISSN 1040-6042.

Horstmann:1992:OCL

[Hor92a] C. Horstmann. The object-based class library in Borland C++. C++ Report, 4(1):15–18, January 1992. CODEN CRPTE7. ISSN 1040-6042.

Horstmann:1992:MCC

[Hor92b] G. Horstmann. The Microsoft C/ C++ Version 7 compiler: A first impression. C++ Report, 4(4):53–57, May 1992. CODEN CRPTE7. ISSN 1040-6042.

Horstmann:1994:EIL

[Hor94] C. S. Horstmann. Extending the iostream library. C++ Report, 6 (4):22–29, May 1994. CODEN CRPTE7. ISSN 1040-6042.

Horstmann:1995:ODT

[Hor95] C. S. Horstmann. OO design tools with C++ code generation. C++ Report, 7(1):56–58, 60–64, January 1995. CODEN CRPTE7. ISSN 1040-6042.

Horstmann:1996:CCL

[Hor96a] C. S. Horstmann. C++ class libraries for numerical programming. C++ Report, 8(1):61–64, 66, January 1996. CODEN CRPTE7. ISSN 1040-6042.

Horstmann:1996:JPS

[Hor96b] C. S. Horstmann. Java programming with Sun’s Java development kit. C++ Report, 8(3):63–70, March 1996. CODEN CRPTE7. ISSN 1040-6042.

Harrison:1996:EPO

[HS96] T. H. Harrison and D. C. Schmidt.

Evaluating the performance of OO network programming toolkits. C++ Report, 8(7):39–46, July 1996. CODEN CRPTE7. ISSN 1040-6042.

Huneke:1991:LCC

[Hun91] I. Huneke. Linking C with C++.

C++ Report, 3(9):12, 16, October 1991. CODEN CRPTE7. ISSN 1040-6042.

Jakubik:1997:PPA

[Jak97] P. Jakubik. Patterns in practice: abstract client. C++ Report, 9(3):38– 44, March 1997. CODEN CRPTE7. ISSN 1040-6042.

Jameson:1996:MPL

[Jam96] K. Jameson. Multiplatform productivity lessons. C++ Report, 8(3): CPS4–CPS6, CPS9, March 1996. CODEN CRPTE7. ISSN 1040-6042.

Jeghers:1996:GDC

[Jeg96] M. Jeghers. GUI development with cross platform toolkits. C++ Report, 8(3):CPS7–CPS9, March 1996. CODEN CRPTE7. ISSN 1040-6042.

Jordan:1995:IOO

[Jor95] D. Jordan. Identifying ODMG objects. C++ Report, 7(3):24, 26– 27, March–April 1995. CODEN CRPTE7. ISSN 1040-6042.

Jordan:1996:OOQ

[Jor96] D. Jordan. OQL: the Object Query Language. C++ Report, 8(2):64– 66, 71–73, February 1996. CODEN CRPTE7. ISSN 1040-6042.

Jain:1997:ECC

[JS97] P. Jain and D. C. Schmidt. Experiences converting a C++ communication software framework to Java. C++ Report, 9(1):50–66, January 1997. CODEN CRPTE7. ISSN 1040-6042.

Kamahele:1996:SCC

[Kam96] M. L. Kamahele. ShMem — a C++ class library for managing objects in shared memory. C++ Report, 8(2): 32–39, 60, February 1996. CODEN CRPTE7. ISSN 1040-6042.

Keffer:1993:DAT

[Kef93] T. Keffer. The design and architecture of Tools.h++. C++ Report, 5(5):28–33, June 1993. CO-DEN CRPTE7. ISSN 1040-6042.

Kontogiorgos:1995:CTA

[KK95] T. Kontogiorgos and M. Kim. A C++ template-based application architecture. C++ Report, 7(2):51– 56, 58, February 1995. CODEN CRPTE7. ISSN 1040-6042.

Kreft:1996:ISC

[KL96a] K. Kreft and A. Langer. Iterators in the Standard C++ Library. C++ Report, 8(10):27–30, 32, November– December 1996. CODEN CRPTE7. ISSN 1040-6042.

Kreft:1996:IST

[KL96b] K. Kreft and A. Langer. Iterators in the Standard Template Library. C++ Report, 8(7):31–37, July 1996. CODEN CRPTE7. ISSN 1040-6042.

Kreft:1997:BIS

[KL97] K. Kreft and A. Langer. Building an iterator for STL and the Standard C++ Library. C++ Report, 9 (2):20–27, February 1997. CODEN CRPTE7. ISSN 1040-6042.

Koenig:1996:CPC

[Koe96a] A. Koenig. Containers and pointer comparisons. C++ Report, 8(10): 23–26, November–December 1996. CODEN CRPTE7. ISSN 1040-6042.

Koenig:1996:CIS

[Koe96b] A. Koenig. Copying and insertion in STL algorithms. C++ Report, 8 (8):65–66, September 1996. CODEN CRPTE7. ISSN 1040-6042.

Koenig:1996:IAS

[Koe96c] A. Koenig. Inheritance, assignment, and slicing. C++ Report, 8(5):71– 72, May 1996. CODEN CRPTE7. ISSN 1040-6042.

Lajoie:1994:EHSb

[Laj94a] J. Lajoie. Exception handling: supporting first-class objects. C++ Report, 6(5):35–39, June 1994. CO-DEN CRPTE7. ISSN 1040-6042.

Lajoie:1994:EHSa

[Laj94b] J. Lajoie. Exception handling: supporting the runtime mechanism. C++ Report, 6(3):36–39, 42– 43, March–April 1994. CODEN CRPTE7. ISSN 1040-6042.

Lajoie:1995:SUT

[Laj95] J. Lajoie. Standards update: type conversions. C++ Report, 7(3):40– 46, March–April 1995. CODEN CRPTE7. ISSN 1040-6042.

Lajoie:1996:CMMb

[Laj96a] J. Lajoie. The C++ memory model: new and delete operators. C++ Report, 8(9):26–33, 44, October 1996. CODEN CRPTE7. ISSN 1040-6042.

Lajoie:1996:CMMa

[Laj96b] J. Lajoie. The C++ memory model. pointer manipulations. C++ Report, 8(8):67–72, September 1996. CODEN CRPTE7. ISSN 1040-6042.

Lajoie:1996:CSC

[Laj96c] J. Lajoie. The C++ Standards Committee’s sixth year-organization and schedule update. C++ Report, 8(9):30–31, October 1996. CODEN CRPTE7. ISSN 1040-6042.

Lajoie:1996:FOR

[Laj96d] J. Lajoie. Function overload resolution: selection of the best viable function. C++ Report, 8(1): 12–17, 56, January 1996. CODEN CRPTE7. ISSN 1040-6042.

Lajoie:1996:MMR

[Laj96e] J. Lajoie. The memory model-type representations. C++ Report, 8(5):

73–76, 81–83, May 1996. CODEN CRPTE7. ISSN 1040-6042.

Lakos:1996:LCSa

[Lak96a] J. Lakos. Large-scale C++ software design. 1. Mentor Graphics experiences using C++ for large systems. C++ Report, 8(6):26–37, June 1996. CODEN CRPTE7. ISSN 1040-6042.

Lakos:1996:LCSc

[Lak96b] J. Lakos. Large-scale C++ software design: levelization, insulation, and packages. 3. avoiding excessive compile- and link-time dependencies. C++ Report, 8(10):42– 56, November–December 1996. CO-DEN CRPTE7. ISSN 1040-6042.

Lakos:1996:LCSb

[Lak96c] J. Lakos. Large-scale C++ software design: physical hierarchy. 2. effective testing, improved maintainability, independent reuse. C++ Report, 8(8):55–64, September 1996. CO-DEN CRPTE7. ISSN 1040-6042.

Lane:1991:DCA

[Lan91] A. Lane. DOS/C++ — application frameworks. C++ Report, 3(7): 14–15, July–August 1991. CODEN CRPTE7. ISSN 1040-6042.

Lea:1993:GCL

[Lea93] D. Lea. The GNU C++ library.

C++ Report, 5(5):24–27, June 1993. CODEN CRPTE7. ISSN 1040-6042.

Leary:1996:CEH

[Lea96] S. Leary. C++ exception handling in multithreaded programs. C++ Report, 8(2):20–31, February 1996. CODEN CRPTE7. ISSN 1040-6042.

Leggett:1992:UCS

[Leg92] B. Leggett. The USL C++ standard components release 2 (end user package). C++ Report, 4(5):69–73, June 1992. CODEN CRPTE7. ISSN 1040-6042.

LeJacq:1991:ECY

[LeJ91] J. P. LeJacq. Education can you do better than C++? C++ Report, 3(10):1, 4–6, November–December 1991. CODEN CRPTE7. ISSN 1040-6042.

Lindner:1996:DSH

[Lin96a] M. P. Lindner. Designing for safe hashing. implementing efficient dynamic search structures. C++ Report, 8(10):36–41, November– December 1996. CODEN CRPTE7. ISSN 1040-6042.

Linenbach:1996:IRB

[Lin96b] T. Linenbach. Implementing reusable binary association in C++. C++ Report, 8(5):40–47, 51, May 1996. CODEN CRPTE7. ISSN 1040-6042.

Lippman:1996:VO

[Lip96] S. Lippman. The varieties of object. C++ Report, 8(6):38–50, June 1996. CODEN CRPTE7. ISSN 1040-6042.

Langer:1997:COO

[LK97] A. Langer and K. Kreft. Combining object-oriented design and generic programming. C++ Report, 9(3): 45–49, 69, March 1997. CODEN CRPTE7. ISSN 1040-6042.

Martin:1992:ACP

[Mar92] R. Martin. Abstract classes and pure virtual functions. C++ Report, 4(6):46–52, July–August 1992. CO-DEN CRPTE7. ISSN 1040-6042.

Martin:1994:PEO

[Mar94] R. Martin. A practical example of object-oriented analysis and design. C++ Report, 6(2):18–23, February 1994. CODEN CRPTE7. ISSN 1040-6042.

Martin:1996:DIP

[Mar96a] R. C. Martin. The Dependency Inversion Principle. C++ Report, 8(6):61–66, June 1996. CODEN CRPTE7. ISSN 1040-6042.

Martin:1996:GOD

[Mar96b] R. C. Martin. Granularity [objectoriented design]. C++ Report, 8(10):57–62, November–December 1996. CODEN CRPTE7. ISSN 1040-6042.

Martin:1996:LSP

[Mar96c] R. C. Martin. The Liskov substitution principle. C++ Report, 8(3):14, 16–17, 20–23, March 1996. CODEN CRPTE7. ISSN 1040-6042.

Martin:1996:OPS

[Mar96d] R. C. Martin. The open-closed principle [software maintenance]. C++ Report, 8(1):37–43, January 1996. CODEN CRPTE7. ISSN 1040-6042.

Martin:1997:DPD

[Mar97a] R. C. Martin. Design patterns for dealing with dual inheritance hierarchies in C++. C++ Report, 9(4):42– 48, April 1997. CODEN CRPTE7. ISSN 1040-6042.

Martin:1997:JCC

[Mar97b] R. C. Martin. Java and C++: a critical comparison. C++ Report, 9 (1):42–49, January 1997. CODEN CRPTE7. ISSN 1040-6042.

Martin:1997:LS

[Mar97c] R. C. Martin. Large-scale stability. C++ Report, 9(2):54–60, February 1997. CODEN CRPTE7. ISSN 1040-6042.

McCluskey:1992:ETI

[McC92] G. McCluskey. An environment for template instantiation. C++ Report, 4(2):1, 4–7, February 1992. CODEN CRPTE7. ISSN 1040-6042.

Meyers:1992:UCE

[Mey92] S. Meyers. Using C++ effectively. approaches to effectiveness. C++ Report, 4(6):34–39, July–August 1992. CODEN CRPTE7. ISSN 1040-6042.

Meyers:1993:UCE

[Mey93] S. Meyers. Using C++ effectively. examining development tools. C++ Report, 5(1):30, 32–35, January 1993. CODEN CRPTE7. ISSN 1040-6042.

Meyers:1995:BOI

[Mey95] S. Meyers. Bounding object instantiations. I. C++ Report, 7(3): 18–22, March–April 1995. CODEN CRPTE7. ISSN 1040-6042.

Meyers:1996:RSP

[Mey96a] S. Meyers. Refinements to smart pointers. C++ Report, 8(10):19– 22, November–December 1996. CO-DEN CRPTE7. ISSN 1040-6042.

Meyers:1996:SPA

[Mey96b] S. Meyers. Smart pointers: automating inheritance and const conversions. C++ Report, 8(8):25–27, September 1996. CODEN CRPTE7. ISSN 1040-6042.

Meyers:1996:SP

[Mey96c] S. Meyers. Smart pointers. I. C++ Report, 8(4):19–23, April 1996. CO-DEN CRPTE7. ISSN 1040-6042.

Meyers:1996:SPN

[Mey96d] S. Meyers. Smart pointers, nullness, and type conversions. C++ Report, 8(6):20–23, June 1996. CO-DEN CRPTE7. ISSN 1040-6042.

Martin:1997:TAP

[MNR97] R. C. Martin, J. W. Newkirk, and B. Rao. Taskmaster: an architecture pattern for GUI applications. C++ Report, 9(3):12–14, 16–23, March 1997. CODEN CRPTE7. ISSN 1040-6042.

Moth:1996:LC

[Mot96] C. Moth. Lightweight classes. C++ Report, 8(5):33–39, May 1996. CO-DEN CRPTE7. ISSN 1040-6042.

Moudgill:1994:RM

[Mou94] M. Moudgill. Reverse-engineering manipulators. C++ Report, 6 (4):30–33, May 1994. CODEN CRPTE7. ISSN 1040-6042.

Mueller:1996:OPL

[Mue96] T. R. Mueller. Overcoming the problems in a large scale C++ project. AT&T’s gamble with C++ as a primary programming language. C++ Report, 8(6):51–52, June 1996. CODEN CRPTE7. ISSN 1040-6042.

Muller:1996:TRH

[Mul96] H. M. Muller. Ten rules for handling exception handling successfully. C++ Report, 8(1):22–36, January 1996. CODEN CRPTE7. ISSN 1040-6042.

Musser:1992:ESC

[Mus92] J. Musser. Extending streambufs: class logstrbuf (OOP). C++ Report, 4(3):51–55, March–April 1992. CO-DEN CRPTE7. ISSN 1040-6042.

Myers:1993:MMC

[Mye93] N. C. Myers. Memory management in C++. I. fine-tuning your programs. C++ Report, 5(6):39– 41, 43, July–August 1993. CODEN CRPTE7. ISSN 1040-6042.

Myers:1995:NUT

[Mye95] N. Myers. A new and useful technique: “traits”. C++ Report, 7 (5):32–35, June 1995. CODEN CRPTE7. ISSN 1040-6042.

Onion:1996:CRC

[OH96a] F. Onion and A. Harrison. Chain of Responsibility and command targets. C++ Report, 8(7):57–59, 62– 63, July 1996. CODEN CRPTE7. ISSN 1040-6042.

Onion:1996:DDE

[OH96b] F. Onion and A. Harrison. Dialog data exchange in MFC. C++ Report, 8(5):48–51, May 1996. CO-DEN CRPTE7. ISSN 1040-6042.

Onion:1996:DDS

[OH96c] F. Onion and A. Harrison. Dynamic data structure serialization in MFC. C++ Report, 8(3):58–62, March 1996. CODEN CRPTE7. ISSN 1040-6042.

Onion:1997:FCF

[OH97] F. Onion and A. Harrison. Framework class factories. C++ Report, 9(1):22–24, January 1997. CODEN CRPTE7. ISSN 1040-6042.

Onion:1996:OPM

[Oni96] F. Onion. Object persistence in MFC. C++ Report, 8(1):57–60, January 1996. CODEN CRPTE7. ISSN 1040-6042.

Onion:1997:MM

[Oni97] F. Onion. Multithreading in MFC.

C++ Report, 9(3):50–53, 56, March 1997. CODEN CRPTE7. ISSN 1040-6042.

Peden:1995:OAA

[Ped95] J. H. Peden. Operator ()() for array of array (of array) class access. C++ Report, 7(5):36–38, 40– 41, June 1995. CODEN CRPTE7. ISSN 1040-6042.

Reed:1993:ETC

[Ree93a] D. Reed. Engineering the transition to C++: a manager’s perspective on realizing the benefits of OOP. C++ Report, 5(1):36, 38–42, January 1993. CODEN CRPTE7. ISSN 1040-6042.

Reed:1993:EPI

[Ree93b] D. R. Reed. Exceptions: pragmatic issues with a new language feature. C++ Report, 5(8):39–44, October 1993. CODEN CRPTE7. ISSN 1040-6042.

Reeves:1996:ESC

[Ree96a] J. Reeves. Exceptions and Standard C++. C++ Report, 8(5):56–59, 62– 64, May 1996. CODEN CRPTE7. ISSN 1040-6042.

Reeves:1996:CE

[Ree96b] J. W. Reeves. Coping with exceptions. C++ Report, 8(3):24–35, March 1996. CODEN CRPTE7. ISSN 1040-6042.

Reeves:1996:EDC

[Ree96c] J. W. Reeves. Exceptions and debugging [in C++]. C++ Report, 8(10):63–71, November–December 1996. CODEN CRPTE7. ISSN 1040-6042.

Reeves:1996:TGE

[Ree96d] J. W. Reeves. Ten guidelines for exception specification. C++ Report, 8(7):64–68, 70, 88, July 1996. CO-DEN CRPTE7. ISSN 1040-6042.

Reeves:1997:RE

[Ree97] J. W. Reeves. Reflections on exceptions. C++ Report, 9(3):57–65, March 1997. CODEN CRPTE7. ISSN 1040-6042.

Reid:1993:CFP

[Rei93] R. J. Reid. C++ as a first programming language. C++ Report, 5(4):41–44, May 1993. CODEN CRPTE7. ISSN 1040-6042.

Remba:1995:SCG

[Rem95] D. Remba. A survey of the computer generated imagery landscape. C++ Report, 7(2):36–41, February 1995. CODEN CRPTE7. ISSN 1040-6042.

Robie:1994:EHO

[Rob94] J. Robie. Event handling in objectoriented databases. C++ Report, 6 (6):24–27, July–August 1994. CO-DEN CRPTE7. ISSN 1040-6042.

Saks:1991:ACU

[Sak91] D. Saks. ANSI C++ update: works in progress. C++ Report, 3(10):13– 14, 16, November–December 1991. CODEN CRPTE7. ISSN 1040-6042.

Saks:1992:ACU

[Sak92] D. Saks. ANSI C++ update: works in progress. C++ Report, 4(1): 11–14, January 1992. CODEN CRPTE7. ISSN 1040-6042.

Sakkinen:1993:HSV

[Sak93] M. Sakkinen. How should virtual bases be initialized (and finalized)? C++ Report, 5(3):44–50, March– April 1993. CODEN CRPTE7. ISSN 1040-6042.

Schmidt:1992:SPC

[Sch92] D. Schmidt. Systems programming with C++ wrappers: encapsulating IPC services with object-oriented interfaces. C++ Report, 4(8):50–54, October 1992. CODEN CRPTE7. ISSN 1040-6042.

Schmidt:1993:CWU

[Sch93a] D. C. Schmidt. A C++ wrapper for UNIX I/O multiplexing: the objectoriented design and implementation of the Reactor. C++ Report, 5(7): 32–43, September 1993. CODEN CRPTE7. ISSN 1040-6042.

Schmidt:1993:EOS

[Sch93b] D. C. Schmidt. Encapsulating operating system IPCs. an objectoriented interface for event-driven UNIX I/O multiplexing. C++ Report, 5(2):43–50, February 1993. CODEN CRPTE7. ISSN 1040-6042.

Schmidt:1994:DAN

[Sch94a] D. Schmidt. A domain analysis of network daemon design dimensions. C++ Report, 6(3):50–59, March– April 1994. CODEN CRPTE7. ISSN 1040-6042.

Schmidt:1994:CSC

[Sch94b] D. C. Schmidt. A case study of C++ design evolution. C++ Report, 6(6): 28–36, July–August 1994. CODEN CRPTE7. ISSN 1040-6042.

Schmidt:1996:CDP

[Sch96] D. C. Schmidt. Connector. A design pattern for actively initializing network services. C++ Report, 8(1):46– 50, 52–56, January 1996. CODEN CRPTE7. ISSN 1040-6042.

Speh:1995:CWW

[SD95] M. Speh and C. Dinnell. C++ and the World Wide Web. C++ Report, 7(1):18–23, January 1995. CODEN CRPTE7. ISSN 1040-6042.

Shur:1997:EIC

[Shu97] Jim Shur. Exploring inner classes in Java 1.1: As Java matures, it begs comparison to the robustness of C++. see how JDK 1.1’s inner classes compare to nested classes. C++ Report, 9(5):18–29, 68, May 1997. CODEN CRPTE7. ISSN 1040-6042.

Sime:1997:GPM

[Sim97] J. Sime. Guarded pointers: moving smart pointers into multithreaded systems. C++ Report, 9(4):32– 41, April 1997. CODEN CRPTE7. ISSN 1040-6042.

Skelly:1992:GHN

[Ske92a] C. Skelly. Getting a handle on the new-handler. C++ Report, 4 (2):16–18, February 1992. CODEN CRPTE7. ISSN 1040-6042.

Skelly:1992:LO

[Ske92b] C. Skelly. In the land of OOP. C++ Report, 4(4):23–26, May 1992. CO-DEN CRPTE7. ISSN 1040-6042.

Stroustrup:1992:RTI

[SL92] B. Stroustrup and D. Lenkov. Runtime type identification for C++. C++ Report, 4(3):32–42, March– April 1992. CODEN CRPTE7. ISSN 1040-6042.

Silva-Lepe:1994:RDC

[SLHS94] I. Silva-Lepe, W. Hursch, and G. Sullivan. A report on Demeter C++. C++ Report, 6(2):24–30, February 1994. CODEN CRPTE7. ISSN 1040-6042.

Soukup:1992:BTC

[Sou92a] J. Soukup. Beyond templates. 1.

(C++). C++ Report, 4(4):27–31, May 1992. CODEN CRPTE7. ISSN 1040-6042.

Soukup:1992:MD

[Sou92b] J. Soukup. Memory-resident databases. C++ Report, 4(2):11–15, February 1992. CODEN CRPTE7. ISSN 1040-6042.

Soukup:1992:SCL

[Sou92c] J. Soukup. Selecting a C++ library.

C++ Report, 4(1):1, 4–6, January 1992. CODEN CRPTE7. ISSN 1040-6042.

Soukup:1996:ACM

[Sou96a] J. Soukup. Another controversy: managing groups of cooperating classes. C++ Report, 8(9):46–47, October 1996. CODEN CRPTE7. ISSN 1040-6042.

Soukup:1996:QP

[Sou96b] J. Soukup. Quality patterns. C++ Report, 8(9):34–44, October 1996. CODEN CRPTE7. ISSN 1040-6042.

Soukup:1997:IP

[Sou97] J. Soukup. Implementing patterns. C++ Report, 9(4):49–50, April 1997. CODEN CRPTE7. ISSN 1040-6042.

Stal:1993:GDA

[SS93] M. Stal and U. Steinmuller. Generic dynamic arrays: integrating dynamic arrays into C++ templates. C++ Report, 5(8):30–35, 38, October 1993. CODEN CRPTE7. ISSN 1040-6042.

Stal:1994:ITE

[SS94] M. Stal and U. Steinmuller. Implementing a text editor using dynamic arrays. C++ Report, 6(4):34–39, 46, May 1994. CODEN CRPTE7. ISSN 1040-6042.

Schmidt:1995:UDP

[SS95] D. C. Schmidt and P. Stephenson. Using design patterns to evolve system software from UNIX to Windows NT. C++ Report, 7(3):47– 60, March–April 1995. CODEN CRPTE7. ISSN 1040-6042.

Stewart:1997:DEH

[SSH97] R. L. Stewart, J. Storey, and D. Huang. Decentralized event handling. C++ Report, 9(2):12–19, February 1997. CODEN CRPTE7. ISSN 1040-6042.

Steinhoff:1991:SV

[Ste91] D. Steinhoff. Saber-C++ version 1.0. C++ Report, 3(6):26–27, 30, June 1991. CODEN CRPTE7. ISSN 1040-6042.

Strickland:1992:ODD

[Str92] H. Strickland. Object databases: a deep look at transparency. C++ Report, 4(5):42–46, June 1992. CO-DEN CRPTE7. ISSN 1040-6042.

Strickland:1993:OOD

[Str93a] H. Strickland. ODMG-93: the object database standard for C++. C++ Report, 5(8):45–48, October 1993. CODEN CRPTE7. ISSN 1040-6042.

Stroustrup:1993:LDU

[Str93b] B. Stroustrup. Library design using C++. C++ Report, 5(5):14–22, June 1993. CODEN CRPTE7. ISSN 1040-6042.

Stroustrup:1993:WCL

[Str93c] B. P. Stroustrup. Why consider language extensions? maintaining a delicate balance. C++ Report, 5 (7):44–51, September 1993. CODEN CRPTE7. ISSN 1040-6042.

Stroustrup:1994:MVF

[Str94] B. Stroustrup. Making a vector fit for a standard. C++ Report, 6 (8):30–34, October 1994. CODEN CRPTE7. ISSN 1040-6042.

Schmidt:1996:CAPb

[SV96a] D. C. Schmidt and S. Vinoski.

Comparing alternative programming techniques for multithreaded CORBA servers. C++ Report, 8 (4):56–66, April 1996. CODEN CRPTE7. ISSN 1040-6042.

Schmidt:1996:CAPc

[SV96b] D. C. Schmidt and S. Vinoski. Comparing alternative programming techniques for multithreaded CORBA servers. C++ Report, 8 (7):47–56, July 1996. CODEN CRPTE7. ISSN 1040-6042.

Schmidt:1996:CAPa

[SV96c] D. C. Schmidt and S. Vinoski. Comparing alternative programming techniques for multithreaded servers. C++ Report, 8(2):50–59, February 1996. CODEN CRPTE7. ISSN 1040-6042.

Schmidt:1997:OES

[SV97] D. C. Schmidt and S. Vinoski. The OMG Events Service. C++ Report, 9(2):37–46, February 1997. CODEN CRPTE7. ISSN 1040-6042.

Tavakkolian:1992:CPP

[Tav92] S. Tavakkolian. Crossing paradigms. A pilgrim’s journey from C to C++. C++ Report, 4(6):40– 45, July–August 1992. CODEN CRPTE7. ISSN 1040-6042.

Thompson:1997:TPC

[TB97] P. Thompson and G. Bumgardner. Threads.h++: a portable C++ library for multithreaded programming. C++ Report, 9(3):24–37, March 1997. CODEN CRPTE7. ISSN 1040-6042.

Teale:1993:TTL

[Tea93] S. Teale. Transplanting a treerecursive LISP algorithm to C++.

C++ Report, 5(4):45–48, May 1993. CODEN CRPTE7. ISSN 1040-6042.

Tichy:1994:GGA

[THN94] W. F. Tichy, J. Heilig, and F. Newbery Paulisch. A generative and generic approach to persistence. C++ Report, 6(1):22–33, January 1994. CODEN CRPTE7. ISSN 1040-6042.

Trinoski:1995:CAD

[TS95] S. Trinoski and D. Schmidt. Comparing alternative distributed programming techniques. C++ Report, 7(4):44–53, May 1995. CO-DEN CRPTE7. ISSN 1040-6042.

VanWyk:1991:SEY

[Van91] C. J. Van Wyk. Simultaneous equations are your friends. C++ Report, 3(9):14–16, October 1991. CODEN CRPTE7. ISSN 1040-6042.

vanLoon:1994:IXW

[van94] R. van Loon. Introducing the X Window System. C++ Report, 6(3): 45–47, March–April 1994. CODEN CRPTE7. ISSN 1040-6042.

VanLoon:1996:PMS

[Van96a] R. Van Loon. The power of multimedia. strengthen your applications to power multimedia features. C++ Report, 8(4):24–28, April 1996. CO-DEN CRPTE7. ISSN 1040-6042.

VanRooyen:1996:ACN

[Van96b] M. Van Rooyen. Alternative C++.

new methods for solving old problems. C++ Report, 8(4):38–42, April 1996. CODEN CRPTE7. ISSN 1040-6042.

Veldhuizen:1995:ET

[Vel95a] T. Veldhuizen. Expression templates. C++ Report, 7(5):26–31, June 1995. CODEN CRPTE7. ISSN 1040-6042.

Veldhuizen:1995:UCT

[Vel95b] T. Veldhuizen. Using C++ template metaprograms. C++ Report, 7(4): 36–38, 40, 42–43, May 1995. CO-DEN CRPTE7. ISSN 1040-6042.

Vilot:1994:IST

[Vil94] M. J. Vilot. An introduction to the Standard Template Library. C++ Report, 6(8):22–29, 35, October 1994. CODEN CRPTE7. ISSN 1040-6042.

Vilot:1995:CSL

[Vil95] M. J. Vilot. The C++ standard library. C++ Report, 7(2):28–34, February 1995. CODEN CRPTE7. ISSN 1040-6042.

Vinoski:1993:DOC

[Vin93] S. Vinoski. Distributed object computing with CORBA. C++ Report, 5(6):32–38, July–August 1993. CO-DEN CRPTE7. ISSN 1040-6042.

Vinoski:1994:MCI

[Vin94] S. Vinoski. Mapping CORBA IDL into C++. C++ Report, 6(7):20– 23, 55, September 1994. CODEN CRPTE7. ISSN 1040-6042.

Vissides:1996:TO

[Vis96] J. Vissides. The trouble with Observer. C++ Report, 8(8):20–24, September 1996. CODEN CRPTE7. ISSN 1040-6042.

Vines:1993:SRT

[VK93] D. Vines and Z. Kishimoto.

Smalltalk’s runtime type support for C++. C++ Report, 5(1):44–52, January 1993. CODEN CRPTE7. ISSN 1040-6042.

Vlissides:1995:PHG

[Vli95] J. Vlissides. Pattern hatchingperspectives from the “gang of four”. C++ Report, 7(3):36– 39, March–April 1995. CODEN CRPTE7. ISSN 1040-6042.

Vlissides:1996:GGS

[Vli96a] J. Vlissides. Generation Gap [software design pattern]. C++ Report, 8(10):12, 14–18, November– December 1996. CODEN CRPTE7. ISSN 1040-6042.

Vlissides:1996:PHP

[Vli96b] J. Vlissides. Protection. I. the Hollywood principle. C++ Report, 8(2): 14, 16–19, February 1996. CODEN CRPTE7. ISSN 1040-6042.

Vlissides:1996:PMP

[Vli96c] J. Vlissides. Protection: membership has its privileges. C++ Report, 8(4):11–13, 16–18, 28, April 1996. CODEN CRPTE7. ISSN 1040-6042.

Vlissides:1996:KS

[Vli96d] J. Vlissides. To kill a singleton.

C++ Report, 8(6):10, 12, 16–19, June 1996. CODEN CRPTE7. ISSN 1040-6042.

Vlissides:1997:TL

[Vli97] J. Vlissides. Type laundering. C++ Report, 9(2):47–52, February 1997. CODEN CRPTE7. ISSN 1040-6042.

Vinoski:1995:DEM

[VS95] S. Vinoski and D. Schmidt. DOC by example: modelling distributed object applications. C++ Report, 7 (2):64–68, February 1995. CODEN CRPTE7. ISSN 1040-6042.

Vinoski:1996:DCD

[VS96] S. Vinoski and D. C. Schmidt. Distributed callbacks and decoupled communication in CORBA. C++ Report, 8(9):48–56, 77, October 1996. CODEN CRPTE7. ISSN 1040-6042.

Webster:1993:CE

[WD93] J. Webster and F. Douglas. C++ is everywhere. C++ Report, 5(4):36– 39, May 1993. CODEN CRPTE7. ISSN 1040-6042.

White:1996:ADU

[Whi96] R. G. White. Advantages and disadvantages of unique representation patterns. C++ Report, 8(8):28–35, September 1996. CODEN CRPTE7. ISSN 1040-6042.

Wilkinson:1992:SCO

[Wil92] N. M. Wilkinson. Subtleties of C++ operator overloading. C++ Report, 4(5):36–41, June 1992. CODEN CRPTE7. ISSN 1040-6042.